OpenStudy (anonymous):

Which of the following rules could represent the function shown in the table? x y -1 -1 0 1 1 3 f(x) = -2x f(x) = 2x - 1 f(x) = 2x + 1

OpenStudy (anonymous):

f(x) = 2x + 1. To find it just put the x values into the formula, 2(-1) +1 = -2+1 = -1 (the first x and y value in the table) do the same for the other questions

OpenStudy (anonymous):

Suck wingspan

OpenStudy (anonymous):

stupid wingspan

OpenStudy (anonymous):

come pinga

OpenStudy (anonymous):

you suck pingaa